OpenStudy (anonymous):

@Catch.me @Mashy @Abhisar Please help me... A particle of mass m travelling with a velocity u makes a head-on elastic cpllision with another particle of mass M which is at rest. If m

OpenStudy (abhisar):

Do you know how to derive this equation \(\huge\sf \boxed{u_2=\frac{m-M}{m+M}u_1}\)

OpenStudy (abhisar):

@MichelleShum ?

OpenStudy (anonymous):

?

OpenStudy (abhisar):

From the above equation we can see that the final velocity v for an object of mass m will be \(\huge\sf \boxed{v=\frac{m-M}{m+M}u}\) Now if m<M, value of v will be negative which means the direction of v will be opposite to u

OpenStudy (anonymous):

ok...Thanks....

OpenStudy (abhisar):

Do you know how to derive the below equation which i used ? \(\huge\sf \boxed{u_2=\frac{m-M}{m+M}u_1}\)

OpenStudy (anonymous):

don't know...

OpenStudy (abhisar):

Have u read ur books carefully ? `Conservation of linear momentum` and `conservation of kinetic energy` in an elastic collision ?

OpenStudy (anonymous):

@Abhisar i know how to get it already...

OpenStudy (abhisar):

That's Great !...in case if u don't u can use this

OpenStudy (anonymous):

ok...